# Diffusion Transformer Captures Spatial-Temporal Dependencies: A Theory for Gaussian Process Data

### v1

@@ C.2.3 Proofs of Other Supporting Lemmas for Proposition 2
$\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}^{(i)}}{\sigma_{t}^{2}}\right\|_{2}^{2}$ $\leq 2\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})\right\|_{2}^{2}+2\mathbb{E}_{\mathbf{z}\sim\mathcal{N}(0,\bm{I})}\|\mathbf{z}/\sigma_{t}\|^{2}$
@@
$\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}^{(i)}}{\sigma_{t}^{2}}\right\|_{2}^{2}\leq 2Nd(R_{s}^{2}+1)\cdot\int_{t_{0}}^{T}\frac{e^{t}{\rm d}t}{e^{t}-1}\leq 2Nd(T+\log(1/t_{0}))(R_{s}^{2}+1).$
@@
$~{}~{}~{}{\left|\mathbb{E}_{\mathbf{v}_{0}}{\left[\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}^{(i)}}{\sigma_{t}^{2}}\right\|_{2}^{2}{\mathbf{1}\left\{\left\lVert\mathbf{v}_{0}\right\rVert\geq R_{0}\right\}}\right]}\right|}$
$\leq 2Nd(R_{s}^{2}+1)(T+\log(1/t_{0}))\cdot\sqrt{\mathbb{P}[\left\lVert\mathbf{v}_{0}\right\rVert^{2}\geq R_{0}^{2}]}$
@@
${\left|\mathbb{E}_{\mathbf{v}_{0}}{\left[\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}^{(i)}}{\sigma_{t}^{2}}\right\|_{2}^{2}{\mathbf{1}\left\{\left\lVert\mathbf{v}_{0}\right\rVert\geq R_{0}\right\}}\right]}\right|}\lesssim\varepsilon(T+\log(1/t_{0})),$
@@
$\quad{\left|\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}}\left\|\widehat{\mathbf{s}}^{(1)}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}}{\sigma_{t}^{2}}\right\|_{2}^{2}-\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}^{(2)}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}}{\sigma_{t}}\right\|_{2}^{2}\right|}$
$={\left|\mathbb{E}_{\mathbf{z}\in\mathcal{N}(\mathbf{0},\bm{I})}{\left[\left\|\widehat{\mathbf{s}}^{(1)}_{t}(\alpha_{t}\mathbf{v}_{0}+\sigma_{t}\mathbf{z})+\frac{\mathbf{z}}{\sigma_{t}}\right\|_{2}^{2}-\left\|\widehat{\mathbf{s}}^{(2)}_{t}(\alpha_{t}\mathbf{v}_{0}+\sigma_{t}\mathbf{z})+\frac{\mathbf{z}}{\sigma_{t}}\right\|_{2}^{2}\right]}\right|}$
@@
Thus, they both have sub-linear growth w.r.t. $\left\lVert\mathbf{z}\right\rVert_{2}$. Now we decompose $\mathbb{E}_{\mathbf{z}\sim\mathcal{N}(\mathbf{0},\bm{I})}{\left[\left\lVert\bm{\phi}_{1}(\mathbf{z})\right\rVert_{2}\left\lVert\bm{\phi}_{2}(\mathbf{z})\right\rVert_{2}\right]}$ by
@@
$\leq\sqrt{(2R_{s}^{2}+2)Nd\Pr{\left[\left\lVert\mathbf{z}\right\rVert_{2}\geq\sigma_{t}^{-1}(R_{1}-\alpha_{t}R_{0})\right]}}.$
@@ C.3 Covering Number of the Multi-layer Transformers
Appendix D Experimental Details in Section 6
@@
From Figure  6, we can see that the weights of the query-key matrix $\mathbf{Q}^{\top}\mathbf{K}$ emphasize the time embedding part of the input across different layers. Interestingly, Figure 7 shows that the weight emphasis on the value matrix $\mathbf{V}$ is on the data patch part instead. This observation aligns with our theoretical construction of the score function using the DiT structure. Specifically, the score matrix (determined by the query-key matrix) captures the kernel $\bm{\Gamma}$, representing temporal dependency, while the value matrix determines the covariance $\bm{\Sigma}$, representing spatial dependency.

### v2

@@ C.2.3 Proofs of Other Supporting Lemmas for Proposition 2
$\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}^{(i)}}{\sigma_{t}^{2}}\right\|^{2}_{2}$ $\leq 2\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})\right\|^{2}_{2}+2\mathbb{E}_{\mathbf{z}\sim\mathcal{N}(0,\bm{I})}\|\mathbf{z}/\sigma_{t}\|^{2}_{2}$
@@
$\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}^{(i)}}{\sigma_{t}^{2}}\right\|^{2}_{2}\leq 2Nd(R_{s}^{2}+1)\cdot\int_{t_{0}}^{T}\frac{e^{t}{\rm d}t}{e^{t}-1}\leq 2Nd(T+\log(1/t_{0}))(R_{s}^{2}+1).$
@@
$~{}~{}~{}{\left|\mathbb{E}_{\mathbf{v}_{0}}{\left[\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}^{(i)}}{\sigma_{t}^{2}}\right\|^{2}_{2}{\mathbf{1}\left\{\left\lVert\mathbf{v}_{0}\right\rVert\geq R_{0}\right\}}\right]}\right|}$
$\leq 2Nd(R_{s}^{2}+1)(T+\log(1/t_{0}))\cdot\sqrt{\mathbb{P}[\left\lVert\mathbf{v}_{0}\right\rVert^{2}_{2}\geq R_{0}^{2}]}$
@@
${\left|\mathbb{E}_{\mathbf{v}_{0}}{\left[\int_{t_{0}}^{T}{\rm d}t\cdot\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}^{(i)}}{\sigma_{t}^{2}}\right\|^{2}_{2}{\mathbf{1}\left\{\left\lVert\mathbf{v}_{0}\right\rVert_{2}\geq R_{0}\right\}}\right]}\right|}\lesssim\varepsilon(T+\log(1/t_{0})),$
@@
$\quad{\left|\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}}\left\|\widehat{\mathbf{s}}^{(1)}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}}{\sigma_{t}^{2}}\right\|^{2}_{2}-\mathbb{E}_{\mathbf{v}_{t}\mid\mathbf{v}_{0}^{(i)}}\left\|\widehat{\mathbf{s}}^{(2)}_{t}(\mathbf{v}_{t})+\frac{\mathbf{v}_{t}-\alpha_{t}\mathbf{v}_{0}}{\sigma_{t}}\right\|^{2}_{2}\right|}$
$={\left|\mathbb{E}_{\mathbf{z}\in\mathcal{N}(\mathbf{0},\bm{I})}{\left[\left\|\widehat{\mathbf{s}}^{(1)}_{t}(\alpha_{t}\mathbf{v}_{0}+\sigma_{t}\mathbf{z})+\frac{\mathbf{z}}{\sigma_{t}}\right\|^{2}_{2}-\left\|\widehat{\mathbf{s}}^{(2)}_{t}(\alpha_{t}\mathbf{v}_{0}+\sigma_{t}\mathbf{z})+\frac{\mathbf{z}}{\sigma_{t}}\right\|^{2}_{2}\right]}\right|}$
@@
Thus, they both have sub-linear growth with respect to $\left\lVert\mathbf{z}\right\rVert_{2}$. Now we can decompose $\mathbb{E}_{\mathbf{z}\sim\mathcal{N}(\mathbf{0},\bm{I})}{\left[\left\lVert\bm{\phi}_{1}(\mathbf{z})\right\rVert_{2}\left\lVert\bm{\phi}_{2}(\mathbf{z})\right\rVert_{2}\right]}$ by
@@
$\leq\sqrt{(2R_{s}^{2}+2)Nd\cdot\Pr{\left[\left\lVert\mathbf{z}\right\rVert_{2}\geq\sigma_{t}^{-1}(R_{1}-\alpha_{t}R_{0})\right]}}.$
@@ C.3 Covering Number of the Multi-layer Transformers
Appendix D Experimental Details in Section 6 and Additional Experiments
D.1 Experimental Details in Section 6
@@
From Figure  8, we can see that the weights of the query-key matrix $\mathbf{Q}^{\top}\mathbf{K}$ emphasize the time embedding part of the input across different layers. Interestingly, Figure 9 shows that the weight emphasis on the value matrix $\mathbf{V}$ is on the data patch part instead. This observation aligns with our theoretical construction of the score function using the DiT structure. Specifically, the score matrix (determined by the query-key matrix) captures the kernel $\bm{\Gamma}$, representing temporal dependency, while the value matrix determines the covariance $\bm{\Sigma}$, representing spatial dependency.
@@
D.2 Additional Experiments
We provide additional numerical results to support and validate our theory: 1) We present comparisons between diffusion transformers and vanilla diffusion models with UNet, and 2) We vary the size of the diffusion transformers and demonstrate the training dynamic of gradually capturing spatial-temporal dependencies.
Comparison of DiT to diffusion with UNet
To conduct a comparison between DiT and diffusion with UNet, we choose one instance of our Gaussian process data with $d=8$ and $N=128$. The covariance function is Gaussian kernel with $\nu=2$ and $\ell=64$. We collect $n=10000$ sequences for training a DiT with $12$ transformer blocks and a UNet-based diffusion model with $4$ down/up sampling procedures. Each of the down/up sampling in UNet consists of $3$ residual and convolution layers so that the DiT and UNet have approximately the same model size.
Each model is sufficiently trained for $400$ epochs, when the training error has converged. In the testing stage, we collect $10000$ samples generated separately from each model and estimate the spatial-temporal dependencies for comparison. As shown in the first column of Figure 11, both DiT and UNet-based diffusion model capture the decay pattern in the temporal correlation. However, DiT exhibits a much better learning result, matching the ground truth. The temporal correlation of the samples generated by UNet-based diffusion model presents a “piecewise” pattern, not as smooth as the ground truth. We conjecture that it is caused by the size of the filter in convolution layers that prevents UNets from learning complete temporal correlation.
Moreover, DiT exhibits significant strength in learning the spatial correlation. As shown in the remaining columns of Figure 11, DiT successfully captures the spatial correlation between two tokens even sufficiently separated, i.e., the correlation is rather weak. In contrast, UNet-based diffusion struggles in learning spatial correlation. We find clear inconsistent patterns of the estimated spatial correlation in the third row, not to mention that the pattern deviates from the ground truth. This result not only demonstrates the surprising learning and generalization capabilities of DiT in sequential data, but also indicates some advantage of DiT over UNet-based diffusion models.
Performance of DiT with varying network size and sequence length
We study the influence of the sequence length $N$ and the number of transformer blocks $L$ on the performance. We choose the sequence length $N$ and the number of transformer blocks $L$ as follows:
1) we fix $L=16$, while change $N$ in $\{2^{4},2^{5},2^{6},2^{7},2^{8}\}$;
2) we fix $N=128$, while change $L$ in $\{1,2,4,8,16\}$.
We adopt a uniform Gaussian process setting of $d=8$, $\nu=2$ and $\ell=64$ as considered in the experiments shown in Figure 11. After training, we independently generate $10000$ sequences for performance evaluation. The metric is the relative error $\epsilon$ of the estimated sample covariance matrix to its ground truth (see a definition in Appendix D.1), which is plotted in Figure LABEL:fig::error-N-D. As shown in the left panel of the figure, the relative error increases mildly as the sequence length increases, which aligns with our $\sqrt{N}$-dependence in Theorem 2, demonstrating DiT’s strength in handling long sequences. The right panel of the figure shows that the performance of DiT improves as the number of transformer blocks increases, yet at a marginally diminishing speed when $L$ is sufficiently large. Using $8$ transformer blocks suffices for an efficient learning in this case, which approximately verifies our construction of a transformer with $\mathcal{O}(\log N)$ blocks.
Figure 11: Visualization of ground truth spatial-temporal dependencies (the first row), spatial-temporal dependencies of the DiT-generated samples (the second row), and the UNet-generated samples (the third row). We visualize the temporal dependencies in the first column, while visualize the spatial dependencies of three pairs $(i,j)\in\{(10,11),(20,30),(30,50)\}$ of tokens respectively, representing short-horizon, medium-horizon, and long-horizon spatial dependencies. For each pair $(i,j)$, we compute the (empirical) covariance matrix between the $i$-th token and the $j$-token and visualize the matrix in a normalized version $\widehat{\gamma}_{i,j}\times\widehat{\bm{\Sigma}}$, where we set $\|\widehat{\bm{\Sigma}}\|_{\rm F}=\left\lVert{\bm{\Sigma}}\right\rVert_{\rm F}$ for a better comparison with the true spatial covariance $\bm{\Sigma}$.
Training dynamic of DiT
To further examine how DiT captures spatial-temporal dependencies, we present the training dynamic of DiT. As shown in Figure 13, the spatial correlation is captured relatively accurately at an early stage of the training (after the first few epochs), while the temporal correlation is learned slower (after epoch 50). We observe the gradual change in the temporal pattern to match eventually the ground truth. In particular, large temporal dependencies are easier to learn, while weak temporal dependencies require more epochs to learn.
Figure 13: Spatial-temporal correlations of samples generated by DiT trained with different number of epochs. We train DiT with $10000$ samples and use Adam optimizer with a mini-batch size of $128$. We visualize the temporal correlation in the first row, and in the second row, we plot the estimated spatial correlation between the $40$-th token and the $60$-token using generated sequences.